OpenStudy (anonymous):

what number of 32% of 75? which of the following equationscould be used to solve the problem?

OpenStudy (zzr0ck3r):

The first one

OpenStudy (anonymous):

32%(x) = 75 \[\frac{ 32*x }{ 100 } = 75\] \[\frac{ 75*100 }{ 32 } = x\]